OpenStudy (mathlegend):

Write the Equation of a line. perpendicular to y=-2x+6 and passing through (-4,2)

OpenStudy (mathlegend):

When finding the perpendicular doesn't that mean it would be the negative reciprocal?

OpenStudy (anonymous):

yes

OpenStudy (ghazi):

well slope of line perpendicular to y=-2x+6 will be m= 1/2 and equation of line will be y=1/2 x+c..by given point you can find c..just put the values of x and y

OpenStudy (mathlegend):

Wait so is it -1/2 or 1/2?

hartnn (hartnn):

+1/2

OpenStudy (mathlegend):

The 2 is already negative so shouldn't I find the positive reciprocal?

OpenStudy (mathlegend):

ok

OpenStudy (anonymous):

yepp

OpenStudy (ghazi):

what you got for c?

OpenStudy (mathlegend):

Is the answer y=1/2x+4 ?

OpenStudy (ghazi):

yes

OpenStudy (ghazi):

correct

OpenStudy (mathlegend):

Thanks everyone!

OpenStudy (ghazi):

:)